Calculus I, Fall 2023 (Thomas’ Calculus Early Transcendentals 13ed), http://www.math.nthu.edu.tw/˜ wangwc/

Brief solutions to selected problems in homework 01

1. Section 2.2: Solutions, common mistakes and corrections:

Figure 1: Section 2.2, problem 81